### v1

# ZeroMerge: Parameter-Free KV Cache Compression for Memory-Efficient Long-Context LLMs

Abstract
The linear growth of key-value (KV) cache memory and quadratic computational complexity pose significant bottlenecks for large language models (LLMs) in long-context processing. While existing KV cache optimization methods address these challenges through token pruning or feature merging, they often suffer from irreversible information loss or require costly parameter retraining. We propose ZeroMerge, a dynamic zero-shot compression framework that achieves efficient cache management through three key innovations: (1) Fine-grained memory allocation guided by multi-dimensional token importance metrics at head-level granularity, (2) A residual merging mechanism that preserves critical context through compensated attention scoring, and (3) Parameter-free adaptation compatible with diverse LLM architectures without retraining. Comprehensive evaluations across LLaMA-2 model demonstrate that ZeroMerge maintains full-cache performance at 5% compression ratios while doubling inference throughput at 40K token lengths. The method effectively balances memory efficiency, generation quality, and deployment flexibility, advancing practical long-context LLM applications. The code is available at https://github.com/SusCom-Lab/ZeroMerge.
Machine Learning, ICML
1 Introduction
The advancement of the Transformer architecture has brought about revolutionary changes in handling sequence data, with large language models (LLMs) being a prime illustration. LLMs have achieved significant progress, surpassing humans in diverse applications. However, handling long contexts has posed challenges, particularly the linear expansion of the KV cache in proportion to the sequence length. For instance, the deployment of a Llama 2 7B model consumes 26GB of memory. When the token length reaches 32K, the 32G of memory occupied by the KV cache becomes the main memory consumer. This, along with the quadratic computational complexity, has impeded LLM integration and performance, spurring research on optimizing the KV cache for enhanced inference efficiency.
Abundant innovative studies on architecture compression alleviate the shortage of memory (Sarah et al., 2024; Ashkboos et al., 2024). Parameter quantization techniques(Dettmers et al., 2022; Xiao et al., 2023a) further reduce memory footprint by representing KV cache with low-bit precision (e.g., 4-bit integers), though at the cost of numerical accuracy preservation. More efforts focus on the management of the KV cache to directly address this challenge. Some sparse methods(Zhang et al., 2024b, a) reserve partial top-important tokens and eliminate others to mitigate the harm to the original generation quality. These selection criteria are based on the historical contribution in the decoding procedure, spatio-temporal relationship, or data distribution. One-sided and short-sighted pre-pruning strategies inevitably disrupt the distribution of KV embeddings and introduce unrecoverable information loss. Hybrid approaches combining quantization and sparsity (Dettmers et al., 2023) attempt to balance accuracy and compression ratio, yet still require delicate bit allocation strategies. Low rank compression techniques are utilized to correct the systematic drawback in some works (Dong et al., 2024). However, the newly introduced architectures or the retraining requirement of the model parameters constitutes a significant burden. One desirable KV management strategy should 1) strictly control the memory consumption to break the context length limits 2) minimize the generation quality degradation 3) be easily applied to various LLM backbones without extra refinement cost.
Motivated by these goals, we proposed ZeroMerge, a more efficient and comprehensive dynamic zero-shot KV management algorithm.
  1. 1.
    Fine-grained memory allocation. ZeroMerge makes full use of the historical contribution of tokens, spatio-temporal characteristics, and its own data distribution characteristics to conduct fine-grained memory management at the head level, making the method more adaptable to different types of long text tasks. It has achieved superior generation quality compared to the sparse strategy and is comparable to the retrained model on models. The performance of 5% of the KV cache bugdet is still on par with the complete model.
  2. 2.
    Low-cost adaptation and deployment. ZeroMerge does not introduce any model parameters and does not require retraining or model fine-tuning to adapt to different compression ratios or various task types. It is adapted to various mainstream model architectures such as Llama, Falcon, and Mistral.
  3. 3.
    Efficient compression strategy. ZeroMerge adopts a light weight operational overhead of linear operational complexity and achieves more than twice the throughput of the original model at a token length of 50K. We have proved that the cache compression mechanism adopted by ZeroMerge does not drown out the original contribution of those intact tokens as the processing sequence grows.
2 Related work
Figure 1: Comparative analysis of KV cache optimization strategies: (a) Model architecture compression reduces both model parameters and KV cache size via multi-query attention heads; (b) Parameter/data quantization trades numerical precision for cache footprint reduction through low-bit representations; (c) Sparse approximation directly addresses memory-compute growth by evicting low-contribution tokens, albeit inducing attention distribution drift; (d) Token merge mitigates sparse caching errors by learning residuals between full and sparse attention outputs, yet introduces overhead from auxiliary networks or fine-tuning; (e) ZeroMerge (proposed) achieves zero-shot compression competitive with sparse+residual methods without added parameters, while demonstrating superior generalization across task domains and compression ratios.
The phenomenon of emergence in large-scale models has ignited new enthusiasm and provided fresh directions for exploring the nonlinear leaps in reasoning and generative capabilities of LLMs—driven by the expansion of model scale and complexity. The unprecedented parameter sizes and intensive computational demands have endowed LLMs with remarkable reasoning and generative abilities, yet they also pose significant challenges, hindering their cost-effective and large-scale deployment across various scenarios. Numerous efforts have been made to address the issue.
2.1 Model architecture compression
Works such as MiniLLM (Gu et al., 2024) and SliceGPT (Ashkboos et al., 2024) have been specifically devoted to ensure that the performance of sample tasks remains intact for more compact models. MiniLLM focuses on knowledge distillation of large language models, aiming to transfer the knowledge from large and powerful models to smaller ones, thereby reducing the number of parameters in the latter. On the other hand, SliceGPT adopts a different approach by replacing weight matrix with a smaller dense matrix. Through these innovative methods, they both strive to strike a balance between controlling the parameter count and maintaining satisfactory performance on relevant tasks. In the research of attention mechanisms, (Voita et al., 2019) pointed out that the contributions of different attention heads to the overall performance of the model in the multi-head self-attention mechanism are not equal. It also provided a novel pruning method at the attention head level. Similarly, a large number of works related to Multi-Query Attention (MQA) enable different attention “head” to share a set of “keys” and “values”, thus significantly reducing the size of these tensors and the memory bandwidth requirements for incremental decoding (Shazeer, 2019; Ainslie et al., 2023).
2.2 Parameter and data quantization
Quantization techniques have emerged as a crucial approach to address the computational and storage challenges of large language models. GPTQ(Frantar et al., 2022) is a one-shot weight quantization method based on approximate second-order information. It can quantize large GPT models with a significant number of parameters in a relatively short time and with minimal accuracy degradation. This enables models to be executed on a single GPU for generative inference and provides end-to-end inference speedups when using different GPUs. INT-FlashAttention(Chen et al., 2024) is the first INT8 quantization architecture compatible with the forward workflow of FlashAttention. It significantly improves the inference speed of FlashAttention on Ampere GPUs by implementing fully INT8 activations and general matrix-multiplication kernels. It also serves as a general token-level post-training quantization framework compatible with other data formats. They not only reduce the model size but also enhance the inference speed and manage quantization errors, making large language models more accessible and efficient for various applications.
2.3 Sparse approximation
Other studies directly focus on the KV cache, another performance bottleneck induced by long sequence tasks, striving to fundamentally address the issues of storage growth and quadratic computational complexity that accompany the increase in sequence sample length. The universality of sparse attention values (Ji et al., 2021; Deng et al., 2024) has provided a new direction for improving the computational efficiency and reducing the memory footprint of models. (Roy et al., 2021) dynamically routes the input tokens to different attention heads based on their content, enabling each attention head to focus only on the relevant parts, thus enhancing the computational efficiency while maintaining the model’s performance. (Wang et al., 2020) takes advantage of the low rank property and projects it into a low-order matrix, thereby avoiding the attention layer with quadratic computational complexity. Another token aware sparse approach, also known as eviction strategy, employ diverse token importance metrics to address memory bottlenecks. The foundational work by (Xiao et al., 2023b) identifies the attention sink phenomenon, where initial tokens disproportionately attract attention scores regardless of positional distance. Building on this insight, StreamingLLM combines a sliding temporal window with persistent retention of initial token KV states, enabling stable infinite-length generation. Similarly, LM-infinite (Han et al., 2024) introduces Lambda-shaped attention masks to differentially strengthen local (recent tokens) and global (starting tokens) semantic modeling. H2O (Zhang et al., 2024b)proposes Heavy Hitter token detection through cumulative attention score thresholds, dynamically evicting non-critical tokens during decoding. Simultaneously, PyramidKV(Zhang et al., 2024a) reveals layer-wise attention divergence - scattered patterns in lower layers gradually consolidate into instruction-critical focus in upper layers. This hierarchy informs its arithmetic-progressive cache allocation, preserving full instruction tokens while progressively reducing intermediate-layer retention.These methods demonstrate multidimensional approaches to token importance quantification. However, these eviction strategies inevitably induce attention distribution drift and error propagation, as permanently discarding low-salience tokens disrupts the intrinsic attention dynamics required for contextually coherent generation (Dao et al., 2022; Xiao et al., 2023b).
2.4 Cache merging
Token merging, initially pioneered in vision transformers (Bolya et al., 2022; Renggli et al., 2022), offers a viable pathway for KV cache optimization by fusing semantically similar tokens. The ToMe (Bolya et al., 2022) framework introduces training-free token fusion through similarity-based matching, enabling layer-wise token reduction without architectural modifications. In textual tasks, Dynamic Memory Compression (Zhang et al., 2024a) extends this concept by adaptively merging spatially proximate tokens, achieving substantial compression while retaining model efficacy through fine-tuning. Further innovations proposed in LESS (Dong et al., 2024) decouple computational bottlenecks by approximating attention mechanisms, where auxiliary networks compress low-priority tokens into compact representations with minimal impact on attention dynamics. From irreversible pruning to feature-preserving merging, from static heuristics to context-aware fusion policies, learnable mergers impose architectural complexity, demanding additional expertise in network tuning and increased computational resources, which complicates deployment.
3 Methodology
Building upon the conceptual foundation of ZeroMerge outlined earlier, this section delineates the technical framework underpinning its zero-shot compression mechanism. We propose a four-component methodology consisting of adaptive budget allocation, context-sensitive contribution evaluation, residual token merging, and stabilized attention projection. The interplay of these components ensures robust generalization across compression ratios and task domains without auxiliary parameters or fine-tuning, addressing limitations of prior eviction and merging strategies.
3.1 Preliminaries
Consider a transformer-based language model with $L$ layers, where each layer contains multi-head attention mechanisms. At decoding step $T$, let $\mathbf{K}_{T}\in\mathbb{R}^{T\times d}$ and $\mathbf{V}_{T}\in\mathbb{R}^{T\times d}$ denote the Key and Value cache matrices for a target attention head, formed by concatenating token-wise vectors up to the current step:
$\mathbf{K}_{T}=[\mathbf{k}_{1},\mathbf{k}_{2},\ldots,\mathbf{k}_{T}]^{\top},\quad\mathbf{V}_{T}=[\mathbf{v}_{1},\mathbf{v}_{2},\ldots,\mathbf{v}_{T}]^{\top},$ (1)
where $\mathbf{k}_{t},\mathbf{v}_{t}\in\mathbb{R}^{d}$ represent the Key and Value vectors for the $t$-th token, and $d$ is the head dimension. For the query vector $\mathbf{q}_{T}\in\mathbb{R}^{d}$ at position $T$, the scaled dot-product attention computes output $\mathbf{o}_{i}$ via:
$a_{t}^{(T)}=\frac{\exp(\mathbf{q}_{T}^{\top}\mathbf{k}_{t}/\sqrt{d})}{\sum_{s=1}^{T}\exp(\mathbf{q}_{T}^{\top}\mathbf{k}_{s}/\sqrt{d})}.\quad\mathbf{o}^{(T)}=\sum_{t=1}^{T}a_{t}^{(T)}\mathbf{v}_{t}$ (2)
This formulation establishes the baseline for analyzing cache compression effects on attention distribution fidelity.
3.2 Budget Allocation
ZeroMerge strategically compresses the original $T$-length cache into a budget $B\ll T$ through tripartite allocation:
$B=B_{c}+B_{r}+B_{p},$ (3)
where $B_{c}$, $B_{r}$, and $B_{p}$ govern context preservation, residual absorption, and proximity maintenance, respectively. The compressed cache matrices aggregate three complementary components:
$\mathbf{K}_{B}=[\mathbf{K}_{c}\|\mathbf{K}_{r}\|\mathbf{K}_{p}],\quad\mathbf{V}_{B}=[\mathbf{V}_{c}\|\mathbf{V}_{r}\|\mathbf{V}_{p}],$ (4)
[…]
The context component $(\mathbf{K}_{c},\mathbf{V}_{c})$ retains top-$B_{c}$ tokens ranked by contribution scores $\mathbf{s}\in\mathbb{R}^{T}$, which quantify contextual saliency. The proximity component $(\mathbf{K}_{p},\mathbf{V}_{p})$ preserves the latest $B_{p}$ tokens, capturing local attention patterns. The residual component $(\mathbf{K}_{r},\mathbf{V}_{r})$ dynamically merges $B_{r}$ historically evicted tokens through learned transformations. This configuration subsumes eviction-based methods when $B_{r}=0$, while enabling continuous interpolation between sparse and dense representations through budget reallocation
3.3 Contribution Evaluation
The context contribution metric $\mathbf{s}^{(T)}\in\mathbb{R}^{T}$ dynamically quantifies each token’s cumulative influence across decoding steps. For the $t$-th token, its score $s_{t}^{(T)}$ evolves through Geometric decay integration of attention activations:
$s_{t}^{(T)}=\begin{cases}\lambda s_{t}^{(T-1)}+a_{t}^{(T)},&T>0\\ 0,&\text{otherwise}\end{cases},$ (5)
where $\lambda\in[0,1]$ is a momentum factor controlling historical preservation. This recursive formulation 1) attenuates early-stage attention signals through geometric decay, 2) maintains temporal locality by restricting updates to causally valid positions, and 3) enables online computation with $O(1)$ update cost per token.
3.4 Residual Token Merging
The residual component dynamically consolidates evicted tokens through similarity-driven aggregation. When merging a candidate token $(\mathbf{k}_{t},\mathbf{v}_{t})$ into $\mathbf{K}_{r}$, we:
  1. 1.
    Select the most compatible residual slot via maximum dot production:
    $\hat{r}=\mathop{\arg\max}\limits_{r\in\{1,\dots,B_{r}\}}\mathbf{k}_{r}^{\top}\mathbf{k}_{t}$ (6)
  2. 2.
    Update the selected slot through momentum-based fusion:
    $\mathbf{k}_{\hat{r}}\leftarrow\frac{w_{\hat{r}}\mathbf{k}_{\hat{r}}+\mathbf{k}_{t}}{w_{\hat{r}}+1},\quad\mathbf{v}_{\hat{r}}\leftarrow\frac{w_{\hat{r}}\mathbf{v}_{\hat{r}}+\mathbf{v}_{t}}{w_{\hat{r}}+1}$ (7)
  3. 3.
    Increment fusion count: $w_{\hat{r}}\leftarrow w_{\hat{r}}+1$
Figure 2 illustrates the dynamic cache evolution under budget parameters $B_{r}=2$, $B_{c}=4$, and $B_{p}=3$ during sequence length expansion from $T=12$ to $T=17$.
Figure 2: KV cache operation process with $B_{r}=2$, $B_{c}=4$, $B_{p}=3$, and sequence length expansion $T=12\rightarrow 17$: residual merging (2 slots), context preservation (top-4) and proximity maintenance (latest 3).
Algorithm 1 ZeroMerge Online Compression
1:  Input: Budgets $B_{c}$, $B_{r}$, $B_{p}$, decay $\lambda$
2:  Init: $\mathbf{K}_{B}\leftarrow(\emptyset,\emptyset,\emptyset)$, $\mathbf{V}_{B}\leftarrow(\emptyset,\emptyset,\emptyset)$
3:  for decoding step $T=1,2,\ldots$ do
4:     Compute attention scores $a_{T,t}$ via (2)
5:     Update contribution scores $\mathbf{s}^{(T)}$ via (5)
6:     $\mathbf{K}_{p}\leftarrow\mathbf{K}_{p}\cup\{\mathbf{k}_{T}\}$, $\mathbf{V}_{p}\leftarrow\mathbf{V}_{p}\cup\{\mathbf{v}_{T}\}$
7:     if $|\mathbf{K}_{p}|>B_{p}$ then
8:        Evict oldest token $(\mathbf{k}_{\text{old}},\mathbf{v}_{\text{old}})$ from $\mathbf{K}_{p}$, $\mathbf{V}_{p}$
9:        $\mathbf{K}_{c}\leftarrow\mathbf{K}_{c}\cup\{\mathbf{k}_{\text{old}}\}$, $\mathbf{V}_{c}\leftarrow\mathbf{V}_{c}\cup\{\mathbf{v}_{\text{old}}\}$
10:        if $|\mathbf{K}_{c}|>B_{c}$ then
11:           Evict lowest-score token $\hat{c}=\arg\min_{c}s_{c}^{(T)}$
12:           $\mathbf{K}_{r}\leftarrow\mathbf{K}_{r}\cup\{\mathbf{k}_{\hat{c}}\}$, $\mathbf{V}_{r}\leftarrow\mathbf{V}_{r}\cup\{\mathbf{v}_{\hat{c}}\}$
13:           if $|\mathbf{K}_{r}|>B_{r}$ then
14:              Merge via (6)–(7)
15:           end if
16:        end if
17:     end if
18:  end for
3.5 Attention Output Stabilization
Algorithm 1 progressively merges low-saliency tokens through dynamic cache updates. To faithfully reconstruct the original attention distribution from compressed representations, we introduce a compensated attention scoring mechanism. The revised attention computation evolves from Eq. 2 to:
$\hat{a}_{t}^{(T)}$ $=\frac{\exp\left(\mathbf{q}_{T}^{\top}\mathbf{k}_{t}/\sqrt{d}+\alpha\log w_{t}\right)}{\sum_{s=1}^{T}\exp\left(\mathbf{q}_{T}^{\top}\mathbf{k}_{s}/\sqrt{d}+\alpha\log w_{s}\right)},$ (8)
$\mathbf{\hat{o}}^{(T)}$ $=\sum_{t=1}^{T}\hat{a}_{t}^{(T)}\mathbf{v}_{t},$
where $w_{j}$ denotes the fusion count of token $j$ (with $w_{j}=1$ for uncompressed tokens), and $\alpha\in(0,1]$ controls residual compensation intensity. This formulation addresses two key challenges:
  • Representation Bias Correction: Merged tokens aggregate multiple historical vectors through Eq. 7, creating a mismatch between their key vectors ($\mathbf{k}_{r}$) and the original value distribution. The $\log w_{j}$ term compensates for this representation shift.
  • Attention Mass Conservation: The compensation term preserves the relative attention mass between compressed and uncompressed tokens, preventing suppression of critical context.
Theorem 3.1.
For any query step $T$ and uncompressed token $i\notin B_{r}$, the revised attention score satisfies $\hat{a}_{i}^{(T)}\geq{a}_{i}^{(T)}$, where ${a}_{i}^{(T)}$ is the original attention score from Eq. 2.
Proof.
Let $r\in B_{r}$ be a residual slot merging $\{r_{1},...,r_{w_{r}}\}$ tokens. For compressed tokens, we establish an upper bound on their attention numerator:
$\text{num}(\hat{a}_{r}^{(T)})$ $=\exp\left(\mathbf{q}_{T}^{\top}\mathbf{k}_{r}/\sqrt{d}+\alpha\log w_{r}\right)$ (9)
$\leq w_{r}\exp\left(\mathbf{q}_{T}^{\top}\mathbf{k}_{r}/\sqrt{d}\right)$
$=w_{r}\exp\left(\frac{1}{w_{r}}\sum_{m=1}^{w_{r}}\mathbf{q}_{T}^{\top}\mathbf{k}_{r_{m}}/\sqrt{d}\right)$
$\leq\sum_{m=1}^{w_{r}}\exp\left(\mathbf{q}_{T}^{\top}\mathbf{k}_{r_{m}}/\sqrt{d}\right),$
where the first inequality uses $\alpha\leq 1$, and the second applies Jensen’s inequality to the convex exponential function. For uncompressed tokens ($w_{i}=1$), we derive:
$\hat{a}_{i}^{(T)}$ $=\frac{\exp(\mathbf{q}_{T}^{\top}\mathbf{k}_{i}/\sqrt{d})}{\sum_{t\notin B_{r}}\exp(\mathbf{q}_{T}^{\top}\mathbf{k}_{t}/\sqrt{d})+\sum_{r\in B_{r}}\text{num}(\hat{a}_{r}^{(T)})}$ (10)
$\geq\frac{\exp(\mathbf{q}_{T}^{\top}\mathbf{k}_{i}/\sqrt{d})}{\sum_{t=1}^{T}\exp(\mathbf{q}_{T}^{\top}\mathbf{k}_{t}/\sqrt{d})}={a}_{i}^{(T)}.$
as the denominator contains compressed tokens’ upper-bounded contributions. ∎
This theoretical guarantee ensures uncompressed tokens maintain their relative attention dominance despite cache compression. The compensation mechanism effectively:
  • Preserves attention mass for critical context tokens
  • Prevents over-amplification of merged token scores
4 Experiment
In this chapter, we aim to verify the effectiveness of our proposed method in terms of performance preservation and memory reduction. Firstly, we introduce the experimental setup, including the selected model architectures, the baseline methods for comparison, and parameters such as the datasets used. Secondly, we verify the effects of our method on memory control and throughput. Moreover, we compare the generation quality of different methods in the field of long texts and tasks. Finally, we discuss the trade-off between memory, time, and performance as well as the impact of relevant hyperparameters on the model.
4.1 Experiment setup
The backbone models selected in the experiment are some small-scale open-sourced LLMs, including LLaMA-2 7B, LLaMa3-8B-Instruct, and Mistral-7B-Instruct. The experimental equipment is an NVIDIA A800-80GB GPU. For the dataset, we choose LongBench, which is a comprehensive benchmark dataset specifically designed for evaluating the performance of different methods on tasks that involve long-context inputs. It contains a wide variety of natural language processing tasks with different lengths of text sequences, covering multiple domains such as news, novels, and academic papers, thus providing a rich and diverse set of test cases for our experiments.
The baselines for comparison are different KV cache management methods. Specifically, they include:
  • Full KV (FKV): Caches all keys and values per token in each layer, serving as the vanilla baseline with maximal computational load but theoretically optimal performance.
  • Heavy Hitter Oracle (H2O): Dynamically manages KV cache using cumulative attention thresholds to balance recent tokens and Heavy Hitters (H2). Leverages layer-wise attention score averaging across queries to identify critical tokens, achieving adaptive compression while maintaining task-specific semantics.
  • LESS: Introduces recurrent KV state synthesis through layer-aware dynamic merging. Combines temporal locality exploitation (retaining recent tokens) with spatial compression via parameterized merging operators. Implements attention rectification to compensate for distribution shifts caused by merged states, enabling sublinear cache growth while preserving contextual coherence.
  • ZeroMerge (Ours): Builds upon merging paradigm but eliminates parametric dependencies through heuristic-based 0-shot merging. Strictly enforces preset cache bounds via two-phase compression: 1) Attention sink preservation for initial/recent tokens, 2) Distance-aware similarity clustering for intermediate tokens. Minimizes embedding distortion through revised attention scoring that accounts for merged token positional densities, achieving better generalization across compression ratios than prior learning-based methods.
4.2 Inference efficiency and resource consumption
Figure 3: VRAM Utilization and Decoding Throughput Across Sequence Lengths. ZeroMerge enforces constant memory footprint (43GB) and sustains 9 tokens/sec decoding rate beyond 54K tokens, eliminating OOM via dynamic KV cache compression.
We evaluate the long - context processing efficiency on a single NVIDIA A800 GPU with 80GB VRAM (Video Random Access Memory). The experiment uses the LLaMA2 - 7B model, which has 32 layers, 32 attention heads, and an embedding dimension of 4096, in float32 precision. The benchmark is composed of a synthetic QA task, where the cache tokens range from 20K to 54K. For the baseline measurements, we use full KV caching without compression. In contrast, the ZeroMerge method enforces a token cache budget of 18K.
The original Llama2-7B model (float32) required 25GB VRAM for parameter loading alone. During prefill phase, processing 20K prompt tokens generated 20GB KV cache across 32 layers, totally consuming 56% of total available VRAM (Fig.3). In baseline inference, KV cache grew linearly with sequence length at 1MB/token. By 54K tokens, cache memory dominated the mast of VRAM, triggering OOM errors. Concurrently, decoding throughput degraded from 9 tokens/sec (20K context) to 4 tokens/sec (52K context)—due to the increased computation of attention score.
By enforcing a 18K token cache ceiling through two-phase compression (attention sink preservation + similarity-based merging), ZeroMerge: Reduced prefill-phase KV cache to 18GB (10% reduction); Maintained constant 43GB VRAM usage during decoding (vs. baseline’s 45 to 79 GB linear growth); Sustained stable throughput at 9 tokens/sec across 20K–54K token contexts; Eliminated OOM constraints through dynamic cache compression.
4.3 Workload-Scalable KV Efficiency
We evaluate ZeroMerge against three baseline methods (FullKV, H2O, LESS) on an NVIDIA A800 GPU (80GB VRAM) using LLaMA2-7B (float32) with diverse workloads: sequence lengths (2K–54K tokens), batch sizes (4–16), and document-QA tasks. Each method is tested under identical decoding constraints (temperature=0.0, greedy sampling) to isolate computational overhead. Metrics include throughput (tokens/sec), latency (sec), and OOM thresholds.
Table 1: Workload-Scalable KV Cache Compression: ZeroMerge Outperforms Baselines in Throughput (tokens/sec) and Latency (seconds) Across Sequence Lengths and Batch Sizes.
Seq.length Batch size Model size throughput(tokens/S) $\Uparrow$ / latency(S) $\Downarrow$
FullKV H2O (5%) Less (5%) ZeroMerge (5%)
4096+4096 4 7B 62.4 / 262.1 43.9 / 372.5 15.0 / 1086.2 77.0 / 212.7
4096+4096 8 7B 65.5 / 500.3 OOM OOM 146.7 / 223.2
2048+2048 16 7B 133.1 / 246.2 86.1 / 380.1 OOM 281.9 / 178.36
4096+4096 16 13B OOM - - 178.2 / 367.6
As demonstrated in Table 1, ZeroMerge consistently outperforms existing methods across diverse workload configurations. At moderate batch sizes (batch=4, 4K+4K tokens), ZeroMerge achieves 23.4% higher throughput (77.0 vs 62.4 tokens/sec) and 18.9% lower latency (212.7s vs 262.1s) compared to FullKV, while H2O and LESS suffer 29.7% and 75.9% performance degradation respectively. Under high-throughput demands (batch=8, 4K+4K tokens), ZeroMerge successfully processes requests where both H2O and LESS encounter OOM errors, delivering 124% throughput improvement (146.7 vs 65.5 tokens/sec) over FullKV. Most notably, at large batch sizes (batch=16, 2K+2K tokens), ZeroMerge sustains 2.1× higher throughput (281.9 vs 133.1 tokens/sec) with 52.8% latency reduction (116.2s vs 246.2s), showcasing its unique capability to decouple computational load from sequence length. These results validate ZeroMerge’s dual advantage, eliminating memory bottlenecks through strict cache budgeting.
4.4 Inference quality
We evaluate text generation quality on the XSum abstractive summarization dataset (16k news articles with single-sentence summaries) using ROUGE-1/2/L metrics. The study compares three KV cache compression methods across two 7B models (Llama2, Falcon) at 20%, 10%, and 5% cache budgets. H2O balances recent tokens and heavy hitters per its original design, while LESS employs its recommended C4-trained merging network. ZeroMerge operates with $\alpha=0.6$ (influence ratio of residual cache ) and $\lambda=0.98$ (the decay rate of the historical contribution) without task-specific training.
Our evaluation demonstrates that ZeroMerge achieves superior text generation quality across varying compression ratios and model architectures compared to existing KV cache compression methods. When compressing the KV cache to 5% of its original size, ZeroMerge maintains near-original performance on Llama2-7B, even slightly exceeding the uncompressed baseline in ROUGE scores, while H2O and LESS exhibit measurable degradation under the same conditions. This advantage becomes more pronounced at higher compression ratios, where ZeroMerge’s performance preservation significantly outperforms both baselines.
On Falcon-7B, a model with distinct architectural characteristics, ZeroMerge demonstrates stronger generalization than parameter-dependent approaches. While LESS suffers severe quality degradation—likely due to its reliance on C4 dataset training—ZeroMerge retains over half of the baseline performance even at extreme 5% compression. In contrast, H2O struggles to balance recent and heavy-hitter tokens effectively, particularly for Falcon’s attention patterns.
The results highlight two critical trends: First, higher compression ratios amplify ZeroMerge’s relative advantages, as its heuristic merging mechanism better preserves semantically critical tokens compared to purely eviction policies. Second, architecture-agnostic design enables consistent cross-model robustness, avoiding the domain adaptation challenges faced by data-driven methods like LESS.
Table 2: Model-Agnostic Text Generation Quality under KV Cache Compression: ROUGE-1/2/L Scores for Llama2-7B and Falcon-7B Across 20%, 10%, and 5% Cache Budgets
method Llama2 7B Falcon 7B
FULL 30.59 / 11.34 / 25.50 27.06 / 8.79 / 22.39
20% H2O 30.83 / 11.43 / 25.71 24.18 / 7.47 / 20.14
20% LESS 30.47 / 11.23 / 25.44 24.90 / 7.97 / 20.76
20% ZeroMerge 31.62 / 12.40 / 26.42 25.19 / 8.69 / 21.34
10% H2O 30.18 / 11.32 / 25.28 13.03 / 3.16 / 11.02
10% LESS 30.74 / 11.22 / 25.58 8.99 / 2.55 / 7.50
10% ZeroMerge 31.83 / 12.47 / 26.75 20.92 / 6.74 / 17.50
5% H2O 28.92 / 10.81 / 24.35 12.02 / 2.08 / 10.36
5% LESS 29.98 / 11.09 / 25.02 7.75 / 1.15 / 6.76
5% ZeroMerge 30.60 / 11.67 / 25.72 15.04 / 3.29 / 12.73
Acknowledgements
Do not include acknowledgements in the initial version of the paper submitted for blind review.
If a paper is accepted, the final camera-ready version can (and usually should) include acknowledgements. Such acknowledgements should be placed at the end of the section, in an unnumbered section that does not count towards the paper page limit. Typically, this will include thanks to reviewers who gave useful comments, to colleagues who contributed to the ideas, and to funding agencies and corporate sponsors that provided financial support.
Impact Statement
Authors are required to include a statement of the potential broader impact of their work, including its ethical aspects and future societal consequences. This statement should be in an unnumbered section at the end of the paper (co-located with Acknowledgements – the two may appear in either order, but both must be before References), and does not count toward the paper page limit. In many cases, where the ethical impacts and expected societal implications are those that are well established when advancing the field of Machine Learning, substantial discussion is not required, and a simple statement such as the following will suffice:
“This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.”
The above statement can be used verbatim in such cases, but we encourage authors to think about whether there is content which does warrant further discussion, as this statement will be apparent if the paper is later flagged for ethics review.