### v2

# ZSMerge: Zero-Shot KV Cache Compression for Memory-Efficient Long-Context LLMs

Abstract
The linear growth of key-value (KV) cache memory and quadratic computational complexity in attention mechanisms pose significant bottlenecks for large language models (LLMs) in long-context processing. While existing KV cache optimization methods address these challenges through token pruning or feature merging, they often incur irreversible information loss or require costly parameter retraining. To this end, we propose ZSMerge, a dynamic KV cache compression framework designed for efficient cache management, featuring three key operations: (1) fine-grained memory allocation guided by multi-dimensional token importance metrics at head-level granularity, (2) a residual merging mechanism that preserves critical context through compensated attention scoring, and (3) a zero-shot adaptation mechanism compatible with diverse LLM architectures without requiring retraining. ZSMerge significantly enhances memory efficiency and inference speed with negligible performance degradation across LLMs. When applied to LLaMA2-7B, it demonstrates a 20:1 compression ratio for key-value cache retention (reducing memory footprint to 5% of baseline) while sustaining comparable generation quality, coupled with triple throughput gains at extreme 54k-token contexts that eliminate out-of-memory failures. The code is available at https://github.com/SusCom-Lab/ZSMerge.
Machine Learning, ICML
1 Introduction
The advancement of the Transformer architecture has revolutionized sequence data processing, with Large Language Models (LLMs) being a prime illustration. LLMs have achieved remarkable progress, surpassing human capabilities in diverse applications. While multi-turn dialogue and long-context understanding are critical scenarios, handling extended sequences remains challenging due to the linear expansion of the Key-Value (KV) cache (which stores intermediate attention keys and values during generation to avoid redundant computations) in proportion to text length. For instance, deploying a LLaMA2-7B model will consume about 26GB of GPU memory; when the token length reaches 32K, the KV cache alone occupies 32G of memory, becoming the primary memory consumption. This, along with the quadratic computational complexity, significantly restricts LLM integration and performance, thus motivating recent research efforts toward optimizing KV cache utilization for enhanced inference efficiency.
To mitigate the large memory footprint of the KV cache, various optimization strategies have emerged. Some modify the core model architecture (Sarah et al., 2024; Ashkboos et al., 2024) or employ quantization techniques for lower precision representation (Dettmers et al., 2022; Xiao et al., 2023a). Other methods directly target the KV cache bottleneck during inference using context-aware techniques. Common among these is sparse approximation or token eviction, retaining only crucial tokens while discarding others (Zhang et al., 2024b, a); however, this can disrupt KV embeddings and cause information loss. To mitigate this, alternative context-aware techniques merge or fuse information from similar tokens (Dong et al., 2024), often using low-rank compression, though these may introduce architectural overhead or require model retraining.
Ideally, an effective and practical KV cache management strategy should 1) strictly control the memory consumption to break the context length limits, 2) minimize the generation quality degradation, and 3) readily applicable across diverse LLM backbones without necessitating fine-tuning or architectural changes (zero-shot compatibility).
Motivated by these goals, we propose ZSMerge, an efficient, comprehensive, and dynamic zero-shot KV management algorithm, with the following key features:
  • Fine-grained memory allocation. ZSMerge makes full use of the historical contribution of tokens, spatio-temporal characteristics, and its own data distribution characteristics to conduct fine-grained, head-level memory management. This adaptability allows it to maintain superior generation quality compared to the sparse strategies and achieve performance comparable to retraining-dependent approaches, even when operating with significantly reduced memory budgets. This ensures effectiveness across diverse long-context tasks.
  • Zero-shot and low-cost Integration. Designed for ease of use, ZSMerge does not introduce any model parameters and thus requires no retraining or model fine-tuning to adapt to different compression ratios or various task types. This also makes it compatible to various mainstream LLM architectures such as LLaMA, Falcon, and Mistral, thereby minimizing deployment overhead.
  • Efficient and throughput-boosting strategy. Characterized by linear operational complexity, ZSMerge’s compression mechanism is computationally lightweight. This efficiency brings significant throughput improvement, e.g., it helps achieve over triple the inference throughput compared to the original model at a token length of 54K. Furthermore, we have proved that the KV cache compressed by ZSMerge can effectively preserves the informational contribution of retained tokens, preventing signal degradation even as context length increases substantially.
2 Related Work
Figure 1: Comparative analysis of KV cache optimization strategies: (A) Sparse approximation directly addresses memory-compute growth by evicting low-contribution tokens, albeit inducing attention distribution drift. (B) Token merge mitigates sparse caching errors, often by learning residuals implicitly or explicitly, yet introduces overhead from auxiliary networks or fine-tuning. (C) ZSMerge (proposed) achieves zero-shot compression competitive with methods requiring learned components, doing so without added parameters while demonstrating superior generalization.
The escalating memory and computational demands of large language models (LLMs), particularly from KV cache growth in long-context scenarios, have spurred diverse optimization strategies. These are broadly categorized into context-agnostic and contextually adaptive approaches.
2.1 Context-Agnostic Optimization
Context-agnostic methods universally reduce resource consumption by altering the model’s structure or numerical representation, irrespective of the input sequence. Foundational techniques include Structural Compression via knowledge distillation (e.g., MiniLLM (Gu et al., 2024)) or matrix factorization (e.g., SliceGPT (Ashkboos et al., 2024)). Attention Mechanism Optimization, such as head pruning (Voita et al., 2019) and key-value sharing in Multi-Query (MQA (Shazeer, 2019)) and Grouped-Query Attention (GQA (Ainslie et al., 2023)), significantly cuts memory bandwidth. Simultaneously, Numerical Optimization tackles precision redundancy through techniques like post-training quantization (e.g., GPTQ (Frantar et al., 2022)), hardware-aligned quantization (e.g., INT-FlashAttention (Chen et al., 2024)), and exploiting low-rank properties (e.g., Linformer (Wang et al., 2020)). These strategies collectively alleviate memory bottlenecks via architectural simplification and precision reduction.
2.2 Contextually Adaptive Optimization
While context-agnostic methods offer baseline efficiency, recent research increasingly aims to directly target the KV cache—a critical bottleneck in long-sequence tasks driven by storage growth and quadratic computational complexity. The observed sparsity in attention values (Ji et al., 2021; Deng et al., 2024) motivates dynamic, context-aware strategies. Figure 1 illustrates the main categories of these adaptive approaches:
(A) Sparse Approximation (Eviction): This class directly addresses memory-compute growth by evicting low-contribution tokens, based on the premise that not all history is equally vital. Work like (Xiao et al., 2023b) identified ”attention sinks,” leading to methods like StreamingLLM (Xiao et al., 2023b) which retains initial tokens alongside a sliding window. LM-infinite (Han et al., 2024) uses specialized masks to balance local and global context. H2O (Zhang et al., 2024b) employs attention score thresholds for dynamic eviction, while PyramidKV (Zhang et al., 2024a) uses layer-wise attention patterns for progressive caching. Dynamic routing (Roy et al., 2021) also fits here. However, eviction strategies inevitably induce attention distribution drift and error propagation by permanently discarding tokens (Dao et al., 2022; Xiao et al., 2023b).
(B) Token Merging: To mitigate information loss from eviction, token merging fuses similar tokens instead of discarding them. Pioneered in vision (e.g., ToMe (Bolya et al., 2022)), it applies to text as well. Dynamic Memory Compression adaptively merges spatially proximate tokens, often requiring fine-tuning. LESS (Dong et al., 2024) uses auxiliary networks to compress tokens with minimal attention impact. While merging mitigates sparse caching errors by learning residuals (implicitly or explicitly), it often introduces overhead from auxiliary networks or fine-tuning, complicating deployment.
(C) ZSMerge (Proposed): Addressing the limitations of both prior approaches, we propose ZSMerge. It achieves zero-shot compression competitive with sparse+residual methods without added parameters, while demonstrating superior generalization across task domains and compression ratios. ZSMerge offers a parameter-free, tuning-free approach, balancing efficiency and generation performance.
3 ZSMerge Methodology
Building upon the conceptual foundation of ZSMerge outlined earlier, this section delineates the technical framework underpinning its zero-shot compression mechanism. We propose a four-component methodology consisting of adaptive budget allocation, context-sensitive contribution evaluation, residual token merging, and stabilized attention projection. The interplay of these components ensures robust generalization across compression ratios and task domains without auxiliary parameters or fine-tuning, addressing limitations of prior eviction and merging strategies.
3.1 Preliminaries
Consider an $L$-layer transformer with multi-head attention mechanisms. For a target attention head at decoding step $T$, the cached Key and Value matrices are defined as:
$\mathbf{K}_{T}=[\mathbf{k}_{1},\mathbf{k}_{2},\ldots,\mathbf{k}_{T}]^{\top}\in\mathbb{R}^{T\times d},$ (1)
$\quad\mathbf{V}_{T}=[\mathbf{v}_{1},\mathbf{v}_{2},\ldots,\mathbf{v}_{T}]^{\top}\in\mathbb{R}^{T\times d},$
where $\mathbf{k}_{t},\mathbf{v}_{t}\in\mathbb{R}^{d}$ represent the Key/Value vectors for the $t$-th token. For the query vector $\mathbf{q}_{T}\in\mathbb{R}^{d}$ at position $T$, the scaled dot-product attention computes output $\mathbf{o}^{(T)}$ via:
$a_{t}^{(T)}$ $=\frac{\exp(\mathbf{q}_{T}^{\top}\mathbf{k}_{t}/\sqrt{d})}{\sum_{s=1}^{T}\exp(\mathbf{q}_{T}^{\top}\mathbf{k}_{s}/\sqrt{d})},$ (2)
$\quad\mathbf{o}^{(T)}$ $=\sum_{t=1}^{T}a_{t}^{(T)}\mathbf{v}_{t}$
This formulation establishes the baseline for analyzing cache compression effects on attention distribution fidelity.
3.2 Budget Allocation
ZSMerge strategically compresses the original $T$-length cache into a budget $B\ll T$ through tripartite allocation:
$B=B_{p}+B_{c}+B_{r},$ (3)
where $B_{p}$, $B_{c}$, and $B_{r}$ govern context preservation, residual absorption, and proximity maintenance, respectively. The compressed cache matrices aggregate three complementary components:
$\mathbf{K}_{B}=[\mathbf{K}_{p}\|\mathbf{K}_{c}\|\mathbf{K}_{r}],\quad\mathbf{V}_{B}=[\mathbf{V}_{p}\|\mathbf{V}_{c}\|\mathbf{V}_{r}],$ (4)
[…]
The proximity component $(\mathbf{K}_{p},\mathbf{V}_{p})$ preserves the latest $B_{p}$ tokens, capturing local attention patterns. The context component $(\mathbf{K}_{c},\mathbf{V}_{c})$ retains top-$B_{c}$ tokens ranked by contribution scores $\mathbf{s}^{(T)}\in\mathbb{R}^{T}$, which quantify contextual saliency. The residual component $(\mathbf{K}_{r},\mathbf{V}_{r})$ dynamically merges $B_{r}$ historically evicted tokens through learned transformations. This configuration subsumes eviction-based methods when $B_{r}=0$, while enabling continuous interpolation between sparse and dense representations through budget reallocation.
3.3 Contribution Evaluation
The contribution scores $\mathbf{s}^{(T)}\in\mathbb{R}^{T}$ dynamically quantify each token’s cumulative influence across decoding steps. For the $t$-th token, its score $s_{t}^{(T)}$ evolves through exponential decay integration of attention activations:
$s_{t}^{(T)}=\begin{cases}\lambda s_{t}^{(T-1)}+a_{t}^{(T)},&T>0\\ 0,&\text{otherwise}\end{cases},$ (5)
where the decay factor $\lambda\in[0,1]$ operates as a temporal discounting parameter analogous to reinforcement learning credit assignment, governing the exponential decay of historical attention contributions. We recommend setting it to a relatively large value to balance the contributions of long-term and short-term attention scores. Throughout this paper, $\lambda$ is configured as $0.98$ unless otherwise specified, as scores from 300 steps earlier become attenuated to approximately 0.002 through exponential decay ($0.98^{300}\approx 0.002$).
3.4 Residual Token Merging
The residual component dynamically consolidates evicted tokens through similarity-driven aggregation. When merging a candidate token $(\mathbf{k}_{t},\mathbf{v}_{t})$ into $\mathbf{K}_{r}$, we adopt the following three-step procedure:
  1. 1.
    Select the most compatible residual slot via maximum dot production:
    $\hat{r}=\mathop{\arg\max}\limits_{r\in\{1,\dots,B_{r}\}}\mathbf{k}_{r}^{\top}\mathbf{k}_{t}$ (6)
  2. 2.
    Update the selected slot through momentum-based fusion:
    $\mathbf{k}_{\hat{r}}\leftarrow\frac{w_{\hat{r}}\mathbf{k}_{\hat{r}}+\mathbf{k}_{t}}{w_{\hat{r}}+1},\quad\mathbf{v}_{\hat{r}}\leftarrow\frac{w_{\hat{r}}\mathbf{v}_{\hat{r}}+\mathbf{v}_{t}}{w_{\hat{r}}+1}$ (7)
  3. 3.
    Increment fusion count: $w_{\hat{r}}\leftarrow w_{\hat{r}}+1$
Figure 2 illustrates the dynamic cache evolution under budget parameters $B_{r}=2$, $B_{c}=4$, and $B_{p}=3$ during sequence length expansion from $T=12$ to $T=17$.
Figure 2: KV cache operation process with $B_{r}=2$, $B_{c}=4$, $B_{p}=3$, and sequence length expansion $T=12\rightarrow 17$: residual merging (2 slots), context preservation (top-4) and proximity maintenance (latest 3).
Algorithm 1 ZSMerge Online Compression
1:  Input: Budgets $B_{p}$, $B_{c}$, $B_{r}$, decay $\lambda$
2:  Init: $\mathbf{K}_{B}\leftarrow(\emptyset,\emptyset,\emptyset)$, $\mathbf{V}_{B}\leftarrow(\emptyset,\emptyset,\emptyset)$
3:  for decoding step $T=1,2,\ldots$ do
4:     Compute attention scores $a_{T,t}$ via (2)
5:     Update contribution scores $\mathbf{s}^{(T)}$ via (5)
6:     $\mathbf{K}_{p}\leftarrow\mathbf{K}_{p}\cup\{\mathbf{k}_{T}\}$, $\mathbf{V}_{p}\leftarrow\mathbf{V}_{p}\cup\{\mathbf{v}_{T}\}$
7:     if $|\mathbf{K}_{p}|>B_{p}$ then
8:        Evict oldest token $(\mathbf{k}_{\text{old}},\mathbf{v}_{\text{old}})$ from $\mathbf{K}_{p}$, $\mathbf{V}_{p}$
9:        $\mathbf{K}_{c}\leftarrow\mathbf{K}_{c}\cup\{\mathbf{k}_{\text{old}}\}$, $\mathbf{V}_{c}\leftarrow\mathbf{V}_{c}\cup\{\mathbf{v}_{\text{old}}\}$
10:        if $|\mathbf{K}_{c}|>B_{c}$ then
11:           Evict lowest-score token $\hat{c}=\arg\min_{c}s_{c}^{(T)}$
12:           if $|\mathbf{K}_{r}|+1\leq B_{r}$ then
13:              $\mathbf{K}_{r}\leftarrow\mathbf{K}_{r}\cup\{\mathbf{k}_{\hat{c}}\}$, $\mathbf{V}_{r}\leftarrow\mathbf{V}_{r}\cup\{\mathbf{v}_{\hat{c}}\}$
14:           else
15:              Merge token $(\mathbf{k}_{\hat{c}},\mathbf{v}_{\hat{c}})$ into $\mathbf{K}_{r}$ via (6)-(7)
16:           end if
17:        end if
18:     end if
19:  end for
3.5 Attention Output Stabilization
Algorithm 1 progressively merges low-saliency tokens through dynamic cache updates. To faithfully reconstruct the original attention distribution from compressed representations, we introduce a compensated attention scoring mechanism. The revised attention computation evolves from Eq. 2 to:
$\hat{a}_{t}^{(T)}$ $=\frac{\exp\left(\mathbf{q}_{T}^{\top}\mathbf{k}_{t}/\sqrt{d}+\alpha\log w_{t}\right)}{\sum_{i=1}^{T}\exp\left(\mathbf{q}_{T}^{\top}\mathbf{k}_{i}/\sqrt{d}+\alpha\log w_{i}\right)},$ (8)
$\mathbf{\hat{o}}^{(T)}$ $=\sum_{t=1}^{T}\hat{a}_{t}^{(T)}\mathbf{v}_{t},$
where $w_{i}$ represents the fusion count of token $i$ (with $w_{i}=1$ for uncompressed tokens). The hyperparameter $\alpha\in(0,1]$ regulates residual compensation intensity, with empirical analysis suggesting an optimal range of $\alpha\in[0.6,0.8]$ for balanced performance. Notably, setting $\alpha=0$ reduces the merging algorithm to a degenerate case equivalent to a pure eviction policy, eliminating residual compensation entirely.
With Eq. 8, we address two key challenges:
  • Representation Bias Correction: Merged tokens aggregate multiple historical vectors via Eq. 7, creating a mismatch between their key vectors ($\mathbf{k}_{r}$) and the original value distribution. The $\log w_{j}$ term compensates for this representation shift.
  • Attention Mass Conservation: The compensation term preserves the relative attention mass between compressed and uncompressed tokens, ensuring residual compensation does not suppress critical uncompressed components.
Theorem 3.1.
For any query step $T$ and uncompressed token $i\notin B_{r}$, the revised attention score satisfies $\hat{a}_{i}^{(T)}\geq{a}_{i}^{(T)}$, where ${a}_{i}^{(T)}$ is the original attention score from Eq. 2.
Proof.
Let $r\in B_{r}$ be a residual slot merging $\{r_{1},...,r_{w_{r}}\}$ tokens. For compressed tokens, we establish an upper bound on their attention numerator:
$\text{num}(\hat{a}_{r}^{(T)})$ $=\exp\left(\mathbf{q}_{T}^{\top}\mathbf{k}_{r}/\sqrt{d}+\alpha\log w_{r}\right)$ (9)
$\leq w_{r}\exp\left(\mathbf{q}_{T}^{\top}\mathbf{k}_{r}/\sqrt{d}\right)$
$=w_{r}\exp\left(\frac{1}{w_{r}}\sum_{m=1}^{w_{r}}\mathbf{q}_{T}^{\top}\mathbf{k}_{r_{m}}/\sqrt{d}\right)$
$\leq\sum_{m=1}^{w_{r}}\exp\left(\mathbf{q}_{T}^{\top}\mathbf{k}_{r_{m}}/\sqrt{d}\right),$
where the first inequality uses $\alpha\leq 1$, and the second applies Jensen’s inequality to the convex exponential function. For uncompressed tokens ($w_{i}=1$), we derive:
$\hat{a}_{i}^{(T)}$ $=\frac{\exp(\mathbf{q}_{T}^{\top}\mathbf{k}_{i}/\sqrt{d})}{\sum_{t\notin B_{r}}\exp(\mathbf{q}_{T}^{\top}\mathbf{k}_{t}/\sqrt{d})+\sum_{r\in B_{r}}\text{num}(\hat{a}_{r}^{(T)})}$ (10)
$\geq\frac{\exp(\mathbf{q}_{T}^{\top}\mathbf{k}_{i}/\sqrt{d})}{\sum_{t=1}^{T}\exp(\mathbf{q}_{T}^{\top}\mathbf{k}_{t}/\sqrt{d})}={a}_{i}^{(T)}.$
as the denominator contains compressed tokens’ upper-bounded contributions. ∎
This theoretical guarantee ensures that uncompressed tokens retain their relative dominance in attention allocation despite cache compression, even as the denominator accounts for upper-bounded contributions from compressed tokens. The compensation mechanism effectively preserves attention mass for critical context tokens while preventing over-amplification of merged token scores, thereby maintaining the integrity of the original attention distribution under compression.
4 Experimental Evaluations
In this chapter, we aim to verify the effectiveness of our proposed method in terms of performance preservation and memory reduction. Firstly, we introduce the experimental setup, including the selected model architectures, the baseline methods for comparison, and parameters such as the datasets used. Secondly, we verify the performances of our method on memory control and throughput. Moreover, we compare the generation quality of different methods in the field of long texts and tasks. Finally, we discuss the trade-off between memory, time, and performance as well as the impact of relevant hyperparameters on the model.
4.1 Experiment Setup
The backbone models selected in the experiment are some small-scale open-sourced LLMs, including LLaMA2-7B, LLaMA3-8B-Instruct, and Mistral-7B-Instruct. The experimental equipment is an NVIDIA A800-80GB GPU. For the dataset, we choose LongBench, which is a comprehensive benchmark dataset specifically designed for evaluating the performance of different methods on tasks that involve long-context inputs. It contains a wide variety of natural language processing tasks with different lengths of text sequences, covering multiple domains such as news, novels, and academic papers, thus providing a rich and diverse set of test cases for our experiments.
The baselines for comparison are different KV cache management methods. Specifically, they include:
  • Full KV (FullKV): Caches all keys and values per token in each layer, serving as the vanilla baseline with maximal computational load but theoretically optimal performance.
  • Heavy Hitter Oracle (H2O): Dynamically manages KV cache using cumulative attention thresholds to balance recent tokens and Heavy Hitters (H2). Leverages layer-wise attention score averaging across queries to identify critical tokens, achieving adaptive compression while maintaining task-specific semantics.
  • LESS: Introduces recurrent KV state synthesis through layer-aware dynamic merging. Combines temporal locality exploitation (retaining recent tokens) with spatial compression via parameterized merging operators. Implements attention rectification to compensate for distribution shifts caused by merged states, enabling sublinear cache growth while preserving contextual coherence.
  • ZSMerge (Ours): Builds upon merging paradigm but eliminates parametric dependencies through heuristic-based zero-shot merging. Strictly enforces preset cache bounds via two-phase compression: 1) Attention sink preservation for initial/recent tokens, 2) Distance-aware similarity clustering for intermediate tokens. Minimizes embedding distortion through revised attention scoring that accounts for merged token positional densities, achieving better generalization across compression ratios than prior learning-based methods.
4.2 Inference Efficiency Gain
To demonstrate the benefits of ZSMerge in improving inference efficiency, we conducted two proof-of-concept experiments. The first (Figure 3) compares the performance of full KV caching and ZSMerge under increasing sequence lengths in a specific inference case. The second (Table 1) evaluates ZSMerge against full KV caching and other baseline methods under various workloads, including different sequence lengths, batch sizes, and model sizes. Below, we summarize the memory and throughput improvements achieved by ZSMerge.
Figure 3: VRAM Utilization and Decoding Throughput Across Sequence Lengths. ZSMerge enforces constant memory footprint (43GB) and sustains 9 tokens/sec decoding rate beyond 54K tokens, eliminating out-of-memory (OOM) via dynamic KV cache compression.
Table 1: Workload-Scalable KV Cache Compression: ZSMerge Outperforms Baselines in Throughput (tokens/sec) and Latency (seconds) Across Sequence Lengths and Batch Sizes.
Seq.length Batch size Model size throughput(tokens/S) $\Uparrow$ / latency(S) $\Downarrow$
FullKV H2O (5%) Less (5%) ZSMerge (5%)
4096+4096 4 7B 62.4 / 262.1 43.9 / 372.5 15.0 / 1086.2 77.0 / 212.7
4096+4096 8 7B 65.5 / 500.3 OOM OOM 146.7 / 223.2
2048+2048 16 7B 133.1 / 246.2 86.1 / 380.1 OOM 281.9 / 178.36
4096+4096 16 13B OOM - - 178.2 / 367.6
4.2.1 Memory Reduction
Specific Case Analysis
In the baseline setup, the LLaMA2-7B model required 25GB of VRAM for parameter loading, with an additional 20GB KV cache generated during the prefill phase for 20K tokens. This linear growth in KV cache size, at 1MB per token, led to OOM errors as sequence lengths approached 54K tokens.
ZSMerge, constrained by an 18K token cache budget, reduced KV cache size by 10% during the prefill phase and maintained VRAM usage at a constant 43GB during decoding. This prevented the baseline’s linear memory growth (up to 79GB) and completely eliminated OOM errors, enabling efficient long-context processing.
4.2.2 Throughput Improvement
Micro-Level Analysis
Decoding throughput for the baseline dropped from 9 tokens/sec at 20K tokens to 4 tokens/sec at 54K tokens due to increasing attention computation overhead. ZSMerge, in contrast, maintained a consistent throughput of 9 tokens/sec across the same sequence length range by dynamically merging less relevant tokens while preserving critical attention information.
Macro-Level Analysis
Under diverse workloads, ZSMerge consistently outperformed the baseline and other methods. For moderate batch sizes (batch=4), it achieved 23.4% higher throughput and 18.9% lower latency than full KV caching. At higher batch sizes (batch=8), ZSMerge handled requests that caused OOM errors for other methods, delivering a 124% throughput improvement. For large batch sizes (batch=16), it sustained 2.1× higher throughput and halved latency compared to full KV caching.
In summary, ZSMerge effectively reduces memory usage and maintains high throughput across various workloads, outperforming full KV caching and other baselines. Its dynamic cache compression mechanism ensures scalability and efficient long-context processing without sacrificing performance.
4.3 Inference Quality Loss
We evaluate text generation quality on the XSum abstractive summarization dataset (16k news articles with single-sentence summaries) using ROUGE-1/2/L metrics. The experiment compares three KV cache compression methods across two 7B models (LLaMA2-7B and Falcon) at 20%, 10%, and 5% cache budgets. H2O balances recent tokens and heavy hitters per its original design, while LESS employs its recommended C4-trained merging network. For ZSMerge, the hyperparameters were set to static values based on empirical validation: the residual compensation intensity ratio ($\alpha$) was fixed at $0.6$ to balance performance across tasks, and the historical contribution decay rate ($\lambda$) was set to $0.98$ to ensure stable context propagation while prioritizing recent attention patterns. These values were selected through comparative analysis against other configurations (e.g., $\alpha\in[0.6,0.8]$ and $\lambda=0.98$ showed optimal trade-offs between compression efficiency and generation quality) and required no task-specific training.
Our evaluation demonstrates that ZSMerge achieves superior text generation quality across varying compression ratios and model architectures compared to existing KV cache compression methods. When compressing the KV cache to 5% of its original size, ZSMerge maintains near-original performance on LLaMA2-7B, even slightly exceeding the uncompressed baseline in ROUGE scores, while H2O and LESS exhibit measurable degradation under the same conditions. This advantage becomes more pronounced at higher compression ratios, where ZSMerge’s performance preservation significantly outperforms both baselines.
On Falcon-7B, a model with distinct architectural characteristics, ZSMerge demonstrates stronger generalization than parameter-dependent approaches. While LESS suffers severe quality degradation-likely due to its reliance on C4 dataset training-ZSMerge retains over half of the baseline performance even at extreme 5% compression. In contrast, H2O struggles to balance recent and heavy-hitter tokens effectively, particularly for Falcon’s attention patterns.
The results highlight two critical trends: First, higher compression ratios amplify ZSMerge’s relative advantages, as its heuristic merging mechanism better preserves semantically critical tokens compared to purely eviction policies. Second, architecture-agnostic design enables consistent cross-model robustness, avoiding the domain adaptation challenges faced by data-driven methods like LESS.
Table 2: Model-Agnostic Text Generation Quality under KV Cache Compression: ROUGE-1/2/L Scores for LLaMA2-7B and Falcon-7B Across 20%, 10%, and 5% Cache Budgets
method LLaMA2-7B Falcon-7B
Full KV 30.59 / 11.34 / 25.50 27.06 / 8.79 / 22.39
20% H2O 30.83 / 11.43 / 25.71 24.18 / 7.47 / 20.14
20% LESS 30.47 / 11.23 / 25.44 24.90 / 7.97 / 20.76
20% ZSMerge 31.62 / 12.40 / 26.42 25.19 / 8.69 / 21.34
10% H2O 30.18 / 11.32 / 25.28 13.03 / 3.16 / 11.02
10% LESS 30.74 / 11.22 / 25.58 8.99 / 2.55 / 7.50
10% ZSMerge 31.83 / 12.47 / 26.75 20.92 / 6.74 / 17.50
5% H2O 28.92 / 10.81 / 24.35 12.02 / 2.08 / 10.36
5% LESS 29.98 / 11.09 / 25.02 7.75 / 1.15 / 6.76
5% ZSMerge 30.60 / 11.67 / 25.72 15.04 / 3.29 / 12.73
5 Conclusion
This study addresses the critical challenge of memory and computational inefficiency in long-context language models through ZSMerge, a dynamic KV cache compression framework. By introducing head-level memory allocation, residual merging with compensated attention scoring, and architecture-agnostic adaptation, the method achieves sublinear memory growth while preserving generation quality across diverse workloads. Empirical validation demonstrates that ZSMerge reduces VRAM consumption by 82% at 54K tokens,triples inference efficiency under a 5% compression ratio, and outperforms eviction-based baselines by 12-34% in text quality metrics—all without task-specific training. The framework advances practical deployment of long-context LLMs by fundamentally decoupling memory constraints from sequence length, enabling efficient processing of extended dialogues and documents while maintaining semantic coherence.
Impact Statement
This paper presents work whose goal is to advance the fields of machine learning and sustainable artificial intelligence. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.